# Real plane separating $(M-2)$-curves of degree ${d}$ and totally real pencils of degree $({d}-3)$

### v1

@@ Abstract.
It is well known that a non-singular real plane projective curve of degree five with five connected components is separating if and only if its ovals are in non-convex position. In this article, this property is set into a different context and generalised to all real plane separating $(M-2)$-curves.
@@ 1. Introduction
Let $C$ be any smooth complex compact algebraic curve equipped with an anti-holomorphic involution $\sigma:C\rightarrow C$, i.e. a smooth real algebraic compact curve. If the real points $C(\mathbb{R})$ of $C$ separates its complex points $C(\mathbb{C})$, i.e. $C(\mathbb{C})\setminus C(\mathbb{R})$ is disconnected, we say that $C$ is of type I or separating. By Harnack-Klein’s inequality [Har76, Kle73], the number $l$ of connected components of $C(\mathbb{R})$ is bounded by the genus $g$ of $C$ plus one . For any fixed $10\leq i\leq g+$ 1, if $l$ equals $g+1-i$, we say that $C$ is an $(M-i)$-curve. The number $l$ is related to the property of separateness of the curve. For example, if $C$ is separating, then $l$ has the parity of $g+1$. Or, if $C$ is an $M$-curve, then $C$ is separating. In this article, we work with non-singular real algebraic plane projective separating $(M-2)$-curves.
@@
The study of smooth real curves of type I and their separating morphisms has been carried out mainly from two points of view: on the one hand from that of abstract curves; on the other hand from that of curves embedded in some ambient surface; e.g. [Hui01], [Gab06],[CH13], [Cop13], [Cop14], [KS20b], [Ore21]. Any real $M$-curve of genus $g$ admits a separating morphism of degree $g+1$, because of Riemann-Roch theorem. In general, for some fixed integers $i,k$ such that $11\leq i\leq g+$ 1 and $k\geq g+1-i$, given a real separating $(M-i)$-curve $C$, it is not evident, a priori, how to construct a separating morphism $f:C\rightarrow\mathbb{P}^{1}_{\mathbb{C}}$ of degree $k$.
On the other hand, [Gab06, Theorem 7.1] states that a genus $g$ real separating curve with $l$ real connected components admits a separating morphism of degree at most $\frac{g+l+1}{2}$.
@@ 2. Proofs and examples
For any fixed $p\in\mathbb{P}^{1}(\mathbb{R})$, pick any configuration $\mathcal{P}$ of $g-2$ distinct points $p_{j_{1}},\dots,p_{j_{g-2}}$ belonging to $f^{-1}(p)$. Applying the notation in Theorem 2.2, take $D_{1}$ as the curve containing $\mathcal{P}$ and an additional real point $q$, different from $p_{j_{g-1}}$. Remark that $D_{1}$ must be reduced; otherwise $D_{1}$ can be written as a product of two real curves $A^{2}B$, where $A$ and $B$ have degree respectively $s$ and $d-3-2s$ with $1\leq s\leq\lfloor\frac{d-3}{2}\rfloor$. But, this leads to a contradiction. In fact, because of the choice of the $g-2$ points, the curve $C$ must intersect $A\cup B$ in at least $2(g-2)$ points, which is not possible by Bézout theorem.
@@
Since $\text{sepgon}(C)=g$, there exists a separating morphism $f:C\rightarrow\mathbb{P}^{1}$ of degree $g$. Therefore for any fixed $p\in\mathbb{P}^{1}(\mathbb{R})$, the points $p_{1},\dots,p_{g-2}$ of $f^{-1}(p)$ belongs to distinct connected components $C_{i}$’s of $C(\mathbb{R})$, where $21\leq i\leq g-$ 2, and the remaining points $p_{g-1},p_{g}$ belongs to the same connected component $C_{g-1}$.
@@
[…]

### v2

@@ Abstract.
It is a well known that a non-singular real plane projective curve of degree five with five connected components is separating if and only if its ovals are in non-convex position. In this article, this property is set into a different context and generalised to all real plane separating $(M-2)$-curves.
@@ 1. Introduction
Let $C$ be any smooth complex compact algebraic curve equipped with an anti-holomorphic involution $\sigma:C\rightarrow C$, i.e. a smooth real algebraic compact curve. If the real points $C(\mathbb{R})$ of $C$ separates its complex points $C(\mathbb{C})$, i.e. $C(\mathbb{C})\setminus C(\mathbb{R})$ is disconnected, we say that $C$ is of type I or separating.
By Harnack-Klein’s inequality [Har76, Kle73], the number $l$ of connected components of $C(\mathbb{R})$ is bounded by the genus $g$ of $C$ plus one . For any fixed $10\leq i\leq g+$ 1, if $l$ equals $g+1-i$, we say that $C$ is an $(M-i)$-curve. The number $l$ is related to the property of separateness of the curve. For example, if $C$ is separating, then $l$ has the parity of $g+1$. Or, if $C$ is an $M$-curve, then $C$ is separating. In this article, we focus on non-singular real algebraic plane projective separating $(M-2)$-curves; see Theorem 1.6.
@@
The study of smooth real curves of type I and their separating morphisms has been carried out mainly from two points of view: on the one hand from that of abstract curves; on the other hand from that of curves embedded in some ambient surface; e.g. [Hui01], [Gab06],[CH13], [Cop13], [Cop14], [KS20b], [Ore21].
Let $C$ be a real separating curve of genus $g$. If $C$ is an $M$-curve, then it admits a separating morphism of degree $g+1$, because of Riemann-Roch theorem. In general, for some fixed integers $i,k$ such that $11\leq i\leq g+$ 1 and $k\geq g+1-i$, if $C$ is an $(M-i)$-curve, it is not evident, a priori, if it exists and how to construct a separating morphism $f:C\rightarrow\mathbb{P}^{1}_{\mathbb{C}}$ of degree $k$.
A remarkable result, due to Gabard [Gab06, Theorem 7.1], states that a genus $g$ real separating curve with $l$ real connected components admits a separating morphism of degree at most $\frac{g+l+1}{2}$.
@@ 2. Proofs and examples
For any fixed $p\in\mathbb{P}^{1}(\mathbb{R})$, pick any configuration $\mathcal{P}$ of $g-2$ distinct points $p_{j_{1}},\dots,p_{j_{g-2}}$ belonging to $f^{-1}(p)$. Applying the notation in Theorem 2.2, take $D_{1}$ as the degree $(d-3)$ curve containing $\mathcal{P}$ and an additional real point $q$, different from $p_{j_{g-1}}$. Remark that $D_{1}$ must be reduced; otherwise $D_{1}$ can be written as $A^{2}\cup B$, where $A$ and $B$ are two real curves and have degree respectively $s$ and $d-3-2s$. with $1\leq s\leq\lfloor\frac{d-3}{2}\rfloor$. But, this leads to a contradiction. In fact, because of the choice of the $g-2$ points, the curve $C$ must intersect $A\cup B$ in at least $2(g-2)$ points, which is not possible by Bézout theorem.
@@
Since $\text{sepgon}(C)=g$, there exists a separating morphism $f:C\rightarrow\mathbb{P}^{1}$ of degree $g$. Therefore for any fixed $p\in\mathbb{P}^{1}(\mathbb{R})$, the points $p_{1},\dots,p_{g-2}$ of $f^{-1}(p)$ belong to distinct connected components $C_{i}$’s of $C(\mathbb{R})$, where $21\leq i\leq g-$ 2, and the remaining points $p_{g-1},p_{g}$ belong to the same connected component $C_{g-1}$.
@@
For the sake of contradiction, let us suppose that there exists a separating morphism $f:C\rightarrow\mathbb{P}^{1}_{\mathbb{C}}$ of degree $6$ such that $f$ has degree $2$ when restricted to a negative oval of $C(\mathbb{R})$. Then, fix some $p\in\mathbb{P}^{1}(\mathbb{R})$ and apply 2.2 taking as $D_{0}$ the line passing through the two points of $f^{-1}(p)$ belonging to the positive oval and the pseudo-line. Up to a choice of the orientation $\mathfrak{O}$ (double arrows in Fig. 2), we get a contradiction with Theorem 2.2 and therefore such $f$ cannot exist. On the other hand sepgon$(C)=6$. This means that all separating morphisms of $C$ must have odd degree on the three negative ovals.